The square poster has sides measuring 2 feet less than the sides of a square sign. If the distance between their area is 52 square feet find the lengths of the sides of the poster and the sign.
Answer by ikleyn(52781)

.
The square poster has sides measuring 2 feet less than the sides of a square sign. If the distance between their area is 52 square feet
find the lengths of the sides of the poster and the sign.
~~~~~~~~~~~~~~~~~~`

Answer. 12 cm (poster) and 14 cm (sign).

x - y = 2,          (1)
x^2 - y^2 = 52.     (2)

====>

x - y = 2,          (1')
(X+y)*(x-y) = 52.   (2')

====>

x - y = 2,          (1'')
x + y = 26.         (2'')


====>

add (1'') and (2'') to get  x = 14.

Then y = 14 - 2 = 12.

Solved.